Matthew J. Miller

Curriculum Vitae

Contact Apex, NC

Phone: Available Upon Request (please email me)

matt@matthewjmiller.net

www.matthewjmiller.net

Employment 1/2007–present Software Engineer Cisco Systems Research Triangle Park, NC

1/2003–9/2005 Research Assistant Coordinated Science Laboratory, University of Illinois at Urbana-Champaign Urbana, IL

Education Ph.D., Computer Science, December 2006 University of Illinois at Urbana-Champaign Urbana, IL

Thesis: Energy Efficiency and Security for Multihop Wireless Networks

Adviser: Prof. Nitin H. Vaidya

Committee: Prof. Klara Nahrstedt, Prof. Jennifer Hou, and Prof. Indranil Gupta GPA: 3.94/4.0

M.S., Computer Science, December 2003 University of Illinois at Urbana-Champaign Urbana, IL

Thesis: Minimizing Energy Consumption in Sensor Networks Using a Wakeup Radio Adviser: Prof. Nitin H. Vaidya

B.S., Computer Engineering, May 2001 Clemson University Clemson, SC

GPA: 4.0/4.0

Minors: Computer Science and Mathematics

Industry Experience Cisco Systems – IOS developer for 7600 router and SIP-400 linecard platforms.

– Helped implement core dump feature which was a finalist for a Reliability, Availability, and Serviceability (RAS) Award within business unit.

– Sustaining for 7600 IP routing platform code. Consistently resolved the most bugs in the group.

– Involved in troubleshooting several high profile customer cases.

– Reviewer for external proposals for Cisco Research.

– Served multiple times as a Cisco judge for Duke University’s graduate networking project class.

Honors US National Science Foundation Fellowship Awarded in 2001 Acceptance Percentage: 16.2% (903/5560)

ASEE National Defense Science and Engineering Graduate Fellowship Awarded in 2001 Acceptance Percentage: 21.3% (285/1339)

Publications Journal Papers

Conference Papers Submitted Papers

Technical Reports

Research Experience Graduate Work My work focuses on security and energy efficiency in wireless multihop networks, particularly sensor networks. In the security domain, I proposed a key distribution protocol to provide symmetric, pairwise keys that, with high probability, are unknown to eavesdroppers. This was the first work to use the underlying wireless channel diversity to address this problem. In the energy efficiency domain, most power save protocols use static sleeping and listening intervals regardless of the network environment. My work looks at adaptively adjusting these intervals in response to network traffic. Additionally, I have proposed methods of using carrier sensing to further improve the energy efficiency of power save protocols and a lightweight protocol to address the energy-latency tradeoff for broadcast dissemination in sensor networks.

Graduate Class Projects

– Exploring the Energy-Latency Trade-off of Broadcasts in IEEE 802.11 Power Save Networks Joint work with Cigdem Sengul in Fall 2003. Designed, analyzed, simulated, and evaluated (using ns-2) a protocol for power save networks and the impact of its parameters on energy and latency. Selected by the professor as one of the three best two-person projects in the class. Published in IEEE ICDCS 2005.

– Improving Fault Tolerance in AODV Joint work with Jungmin So in Fall 2002. Designed, simulated, and evaluated (using ns-2) techniques to maintain multiple routes in an ad-hoc routing protocol.

– Improving Connectivity in a Scatternet Formation Algorithm Joint work with Cristina L. Abad in Spring 2002. Designed, simulated, and evaluated (using a custom built simulator written in C) a protocol to provide greater connectivity in Bluetooth scatternet formation.

– Log Correlation for Intrusion Detection Group project in Spring 2003. Investigated how information from various system logs can be used to identify specific attacks. Subsequent work by some group members led to a publication based on the project:

– Tools for Middle School Students to Create Vignettes Joint work with Jeffrey Naisbitt and Naomi Caldwell in Spring 2003. Designed, implemented, and did user-testing on a tool (using Java Swing) to allow students to create life stories using an instant messenger-like interface.

Awards

– NSF Student Travel Grant (US $500) for IEEE BROADNETS 2004

– NSF Student Travel Grant (US $750) for ACM SenSys 2004

– DARPA/NSF Student Travel Grant (US $400–750) for IEEE MASS 2005

Undergraduate Work

– Participated in the NSF-funded Summer Undergraduate Research Experience (SURE) at Clemson University in 1999.

– Worked in Parallel Architecture Research Lab (PARL) at Clemson University for two years.

– Projects included designing user interfaces (using Java Swing) for scientific computing problem solving environments to allow message passing between modules and array partitioning.

External Reviewer

– IEEE Transactions on Mobile Computing, IEEE Transactions on Dependable and Secure Computing, IEEE Transactions on Wireless Communications, IEEE Transactions on Information Forensics and Security, and IEEE Communications Magazine.

– MobiQuitous 2004, IEEE WCNC 2004, IEEE ICC 2005, IEEE MASS 2006, and IEEE VTC 2006.

– Cisco Research funding proposals.

– Judge for Duke University Computer Science graduate class projects (2007).

Memberships 1997–present Member, Institute of Electrical and Electronics Engineers (IEEE)

Technical Skills Proficient

C, Python, C++, Java, Perl, LaTeX, Matlab, ns-2 Network Simulator (C++ based), TinyOS/TOSSIM Familiar

Tcl/Tk, ML, Prolog, Lisp, PHP

Extracurricular 2004–2005 Large Group Coordinator UIUC Graduate InterVarsity Christian Fellowship

– Responsible for contacting and scheduling about 20 speakers – Responsible for arranging facility and equipment reservations – Helped restructure the chapter budget

Miscellaneous Citizenship: United States of America Marital Status: Married Erdös Number: ≤ 4

November 1, 2010